OpenStudy (anonymous):

A firefighter jumps out of the window into a net ,if the window is 25m above the net,at what speed does the firefighter hit the net?and how long will it take the firefighter to hit the net?

OpenStudy (anonymous):

G = 9.8 m/s^2. the formula to calculate distance is: \[Dist = (1/2) G t^2\] (see: http://en.wikipedia.org/wiki/Equations_for_a_falling_body) So substituting we get \[25 = (1/2) 9.8 t^2\] \[t^2 = 50/9.8\] \[t = \sqrt{50/9.8}\] \[t = 5\sqrt{2}/\sqrt{9.8}\]